# LiDAR-based End-to-end Temporal Perception for Vehicle-Infrastructure Cooperation

### v1

@@ III-F End-to-End Tracking
The end-to-end tracking capability of LET-VIC is powered by an advanced architecture inspired by the TrackFormer framework [6, 5, 34], which unifies object detection and tracking tasks within a single Transformer-based pipeline. This integration bypasses the conventional two-step tracking-by-detection approach, where detection is followed by post-processing for tracking. Instead, LET-VIC directly associates objects across frames during detection, allowing for joint optimization of both tasks.
@@ IV-A1 Dataset
We use the V2X-Seq-SPD dataset [9], which contains complex traffic intersection scenes captured at 10Hz using both vehicle-side and infrastructure-side sensors. The dataset provides detailed 3D tracking annotations for multiple object classes, such as cars, pedestrians, and cyclists. The dataset is split into train/validation/test (5:2:3).
@@ IV-A2 LET-VIC Implementation Details
We extract V2X-Seq-SPD [9] at a frequency of 10Hz to create the V2X-Seq-SPD-10Hz dataset, aligning with nuScenes [41]. The ego vehicle-side BEV range configuration is $[-51.2,51.2,-51.2,51.2]$ meters with grids sized at 0.512 meters $\times$ 0.512 meters. Considering the infrastructure-side LiDAR’s fan-shaped perception, the infrastructure-side BEV range configuration is $[0,102.4,-51.2,51.2]$ meters with grids sized at 0.512 meters $\times$ 0.512 meters. Our evaluation encompasses all tasks utilizing cooperative-view annotations, and the experiments are conducted using the NVIDIA A100 GPU.
TABLE I: Performance Comparison under Different Delay Conditions on the V2X-Seq-SPD 10Hz Dataset V2VNet, FFNet, and PointPillars follows the tracking-by-detection paradigm with multi-view point cloud feature fusion, LET-V is an end-to-end tracking model with single-view point cloud data input, LET-VIC is an end-to-end tracking model that fuses multi-view point cloud data feature. Detection uses mAP and tracking uses AMOTA and AMOTP.
@@ IV-A3 Baseline Settings
We select single-vehicle temporal perception and non-end-to-end VIC perception as two baselines, comparing our approach against state-of-the-art (SOTA) LiDAR-based VIC perception methods, including V2VNet, FFNet, and PointPillars.
@@
  • V2VNet: Detection is performed using V2VNet [18], paired with AB3DMOT [43] for tracking in a tracking-by-detection framework. The model is trained on fused features from both ego-vehicle and infrastructure point clouds. Evaluation includes delay scenarios of 0 ms, 100 ms, 200 ms, and 300 ms.
  • FFNet: Detection is performed using FFNet [22], paired with AB3DMOT [43] for tracking in a tracking-by-detection framework. The model is trained on fused features from both ego-vehicle and infrastructure point clouds. Evaluation includes delay scenarios of 0 ms, 100 ms, 200 ms, and 300 ms.
  • PointPillars: Detection is performed using PointPillars [38], paired with AB3DMOT [43] for tracking in a tracking-by-detection framework. The model is trained on fused features from both ego-vehicle and infrastructure point clouds. Evaluation includes delay scenarios of 0 ms, 100 ms, 200 ms, and 300 ms.
@@ IV-B Experiment Results
Figure 7: Visualization with V2X-Seq-SPD. The visualization comparison clearly demonstrates that our LET-VIC model achieves superior results compared to other baseline models. Each row represents the perception results of each model at time t. Each column shows the temporal perception results of a specific model, with the same color box used for each tracked object. Panel (a) shows the temporal perception results of LET-VIC, (b) represents LET-V, (c) shows PointPillars, (d) shows FFNet, (e) shows V2VNet, and (f) represents ground truth.
Table I and Fig. 6 report the comparative detection and tracking performance of different methods on the V2X-Seq-SPD 10Hz Dataset [9], and LET-VIC shows superior performance on all metrics. Visualizations of each model’s inference results are shown in the fig. 7. The comparison of LET-VIC and LET-V, both in the benchmark results and visualizations, demonstrates the advantage of VIC over single-vehicle perception, underscoring the value of incorporating infrastructure-side information. Furthermore, the evaluation and visual results of LET-VIC, PointPillars, FFNet, and V2VNet highlight the benefits of end-to-end temporal perception, showing that integrating historical temporal information significantly enhances perception performance.
@@ IV-B1 Comparative Results of VIC Perception and Single-Vehicle Perception
The comparative results of LET-VIC and LET-V demonstrate that LET-VIC consistently outperforms LET-V under various delay conditions, highlighting its strong robustness to latency. Leveraging the VIC Cross-Attention module, LET-VIC effectively integrates BEV features from infrastructure sensors, significantly enhancing detection accuracy and tracking performance with minimal degradation, even under delayed conditions. Specifically, in detection performance, LET-VIC achieves substantial improvements over LET-V, with +15.0% mAP at 0ms, +12.3% mAP at 100ms, +9.8% mAP at 200ms and +7.0% mAP at 300ms. For tracking performance, LET-VIC outperforms LET-V with +17.3% AMOTA at 0 ms, +16.4% AMOTA at 100 ms, +14.6% AMOTA at 200 ms and +10.3% AMOTA at 300 ms. Moreover, LET-VIC consistently achieves lower AMOTP values than LET-V, with -0.413 AMOTP at 0 ms, -0.397 AMOTP at 100 ms, -0.338 AMOTP at 200 ms and -0.302 AMOTP at 300 ms, reflecting superior localization accuracy. Additionally, as shown in Fig. 6, although each metric shows a slight decline within a latency range up to 300 ms, LET-VIC still surpasses LET-V by +7.0% in mAP, +10.3% in AMOTA, and -0.302 in AMOTP at 300 ms delay, demonstrating its ability to effectively fuse infrastructure features under certain delays and enhance perception capabilities, showcasing robustness to latency.
TABLE II: Importance of CEC in LET-VIC on the V2X-Seq-SPD 2Hz Dataset at Different Latency Levels.
@@ IV-B2 Comparative Results of End-to-End Paradigms and Tracking-by-Detection Paradigms
The comparative results between LET-VIC and models such as V2VNet [18], FFNet [22], and PointPillars [38] indicate that the end-to-end temporal perception approach of LET-VIC significantly outperforms traditional tracking-by-detection models across various delay settings. Specifically, LET-VIC achieves substantial improvements in detection performance, with at least +13.7% mAP at 0ms, +12.3% mAP at 100ms, +11.0% mAP at 200ms and +8.5% mAP at 300ms compared to the other models. For tracking performance, LET-VIC outperforms V2VNet, FFNet, and PointPillars with at least +13.1% AMOTA at 0 ms, +12.1% AMOTA at 100 ms, +12.7% AMOTA at 200 ms and +9.2% AMOTA at 300 ms. Moreover, LET-VIC consistently achieves lower AMOTP values, with at least -0.122 AMOTP at 0 ms, -0.117 AMOTP at 100 ms, -0.100 AMOTP at 200 ms and -0.063 AMOTP at 300 ms, reflecting superior localization accuracy. Additionally, Compared to V2VNet, FFNet, and PointPillars, LET-V demonstrates that even with single-vehicle perception, the end-to-end approach significantly improves detection and tracking. Its performance approaches that of PointPillars and surpasses both FFNet and V2VNet. This highlights the effectiveness of the end-to-end paradigm in enhancing perception performance, even without infrastructure data.
IV-C Ablation Study
@@ IV-C2 LET-VIC Implementation Details
Due to limited computational resources, this section of experiments extracts the V2X-Seq-SPD dataset at a reduced frequency of 2Hz to generate the V2X-Seq-SPD-2Hz dataset. Additionally, we reduced the feature dimensionality in the model from 256 dimensions to 64 dimensions to expedite training and evaluation. Other settings remain consistent with those described in Section IV-A2.
@@ IV-C3 Analysis
The results presented in Table II demonstrate that LET-VIC with the CEC module consistently outperforms the standard LET-VIC across all delay settings, highlighting the CEC module’s effectiveness in enhancing perception robustness. Specifically, LET-VIC with calibration error compensation achieves +6.4% mAP and +4.7% AMOTA improvements, along with a -0.038 AMOTP reduction at 0 ms delay; +7.5% mAP, +5.5% AMOTA, and -0.079 AMOTP at 100 ms delay; +5.0% mAP, +1.8% AMOTA, and -0.044 AMOTP at 200 ms delay; and +5.2% mAP, +4.2% AMOTA, and -0.042 AMOTP at 300 ms delay. By compensating for sensor calibration errors, this module improves multi-view feature fusion, leading to better overall performance and demonstrating LET-VIC’s robustness to calibration inaccuracies.

### v2

@@ III-F End-to-End Tracking
The end-to-end tracking capability of LET-VIC, shown in part (d) of Fig. 3, is powered by an advanced architecture inspired by the TrackFormer framework [6, 5, 48], which unifies object detection and tracking tasks within a single Transformer-based pipeline. This integration bypasses the conventional two-step tracking-by-detection approach, where detection is followed by post-processing for tracking. Instead, LET-VIC directly associates objects across frames during detection, allowing for joint optimization of both tasks.
@@ IV-A1 Dataset
We conduct experiments on two datasets: the real-world V2X-Seq-SPD dataset [9] and the simulated V2X-Sim dataset [32], collected from Carla.
V2X-Seq-SPD [9] is a real-world dataset designed for VIC temporal perception. It consists of 15,371 pairs of cooperative frames captured at 10Hz across 95 scenes. Each frame pair contains synchronized sensor data from both vehicle-side and infrastructure-side perspectives. We conduct training on 7,445 frame pairs from 46 scenes and evaluate on 3,316 frame pairs from 21 validation scenes. This dataset is used for comparative experiments and ablation studies, with results reported on the validation set.
V2X-Sim [32] is a simulation dataset collected from virtual environments in Carla. It comprises 10,000 cooperative frames recorded at 5Hz across 100 scenes. To validate the generalizability of our approach, we conduct VIC temporal perception experiments on this dataset. In each scene, we select one road-side unit (RSU) and one CAV to establish cooperative perception conditions. We train on 5,000 frame pairs from 50 scenes and evaluate on 2,000 frame pairs from 20 validation scenes.
@@ IV-A2 LET-VIC Implementation Details
We extract V2X-Seq-SPD [9] at a frequency of 10Hz to create the V2X-Seq-SPD-10Hz dataset [9], aligning with nuScenes [56]. The ego vehicle-side BEV range configuration is $[-51.2,51.2,-51.2,51.2]$ meters with grids sized at 0.512 meters $\times$ 0.512 meters. Considering the infrastructure-side LiDAR’s fan-shaped perception, the infrastructure-side BEV range configuration is $[0,102.4,-51.2,51.2]$ meters with grids sized at 0.512 meters $\times$ 0.512 meters. We also evaluate our method on the V2X-Sim dataset [32], which is extracted at 5Hz. To establish cooperative perception conditions, we select one RSU and one CAV per scene. The ego vehicle-side BEV range configuration remains the same as in V2X-Seq-SPD [9], while the infrastructure-side BEV range configuration is set to $[-51.2,51.2,-51.2,51.2]$ meters, with the same grid resolution of 0.512m $\times$ 0.512m. Our evaluation encompasses all tasks utilizing cooperative-view annotations, and the experiments are conducted using the NVIDIA A100 GPU. To assess the computational feasibility of LET-VIC, we analyze its complexity in terms of floating point operations per second (FLOPs) and parameter count. LET-VIC requires 514.86 G FLOPs and consists of 30.11 M parameters. These values reflect the model’s capability to efficiently leverage vehicle-infrastructure cooperation while maintaining practical computational overhead. The balance between accuracy and efficiency makes LET-VIC well-suited for real-world deployment.
TABLE I: Performance Comparison under Different Delay Conditions on the V2X-Seq-SPD-10Hz Dataset V2VNet [18], FFNet [22], and PointPillars [53] follows the tracking-by-detection paradigm with multi-view point cloud feature fusion, LET-V is an end-to-end tracking model with single-view point cloud data input, LET-VIC is an end-to-end tracking model that fuses multi-view point cloud data feature. Detection uses mAP and tracking uses AMOTA and AMOTP.
@@ IV-A3 Baseline Settings
We select single-vehicle temporal perception and non-end-to-end VIC perception as two baselines, comparing our approach against state-of-the-art (SOTA) LiDAR-based VIC perception methods, including V2VNet [18], FFNet [22], and PointPillars [53].
@@
  • V2VNet [18]: Detection is performed using V2VNet [18], paired with AB3DMOT [58] for tracking in a tracking-by-detection framework. The model is trained on fused features from both ego-vehicle and infrastructure point clouds. Evaluation includes delay scenarios of 0 ms, 100 ms, 200 ms, and 300 ms.
  • FFNet [22]: Detection is performed using FFNet [22], paired with AB3DMOT [58] for tracking in a tracking-by-detection framework. The model is trained on fused features from both ego-vehicle and infrastructure point clouds. Evaluation includes delay scenarios of 0 ms, 100 ms, 200 ms, and 300 ms.
  • PointPillars [53]: Detection is performed using PointPillars [53], paired with AB3DMOT [58] for tracking in a tracking-by-detection framework. The model is trained on fused features from both ego-vehicle and infrastructure point clouds. Evaluation includes delay scenarios of 0 ms, 100 ms, 200 ms, and 300 ms.
TABLE II: Performance Comparison under Ideal Communication Conditions on the V2X-Seq-SPD-10Hz and V2X-Sim 5Hz Datasets.
Model V2X-Seq-SPD V2X-Sim
[…]
LET-V 0.456 0.467 1.113 0.187 0.093 1.734
LET-VIC 0.606 0.640 0.700 0.209 0.160 1.606
@@ IV-B Experiment Results
Figure 7: Visualization with V2X-Seq-SPD. The visualization comparison clearly demonstrates that our LET-VIC model achieves superior results compared to other baseline models. Each row represents the perception results of each model at time $t$. Each column shows the temporal perception results of a specific model, with the same color box used for each tracked object. To further illustrate perception performance, we introduce semi-transparent gray dashed missed-detection boxes to highlight undetected objects by each model. LET-VIC exhibits the fewest missed-detection boxes, demonstrating its superior perception capability. Panel (a) shows the temporal perception results of LET-VIC, (b) represents LET-V, (c) shows PointPillars [53], (d) shows FFNet [22], (e) shows V2VNet [18], and (f) represents ground truth.
Table I and Fig. 6 report the comparative detection and tracking performance of different methods on the V2X-Seq-SPD-10Hz Dataset [9], where LET-VIC shows superior performance across all metrics. The comparison between LET-VIC and LET-V, both in the benchmark results and visualizations, demonstrates the advantage of VIC over single-vehicle perception, underscoring the value of incorporating infrastructure-side information. Table II further validates the generalizability of our method through cross-dataset experiments on the V2X-Sim Dataset [32]. Moreover, the evaluation and visual results of LET-VIC, along with those of PointPillars [53], FFNet [22], and V2VNet [18], highlight the benefits of end-to-end temporal perception, showing that integrating historical temporal information significantly enhances perception performance. The inference results of each model are visualized in Fig. 7. As shown in the figure, LET-VIC demonstrates robust object tracking across frames, as evidenced by the consistent color-coding of tracked objects, while the noticeably fewer semi-transparent gray dashed missed-detection boxes further underscore its enhanced perception capability and reduced detection failures.
@@ IV-B1 Comparative Results of VIC Perception and Single-Vehicle Perception
The comparative results of LET-VIC and LET-V demonstrate that LET-VIC consistently outperforms LET-V under various delay conditions, highlighting its strong robustness to latency. This improvement is primarily attributed to vehicle-infrastructure coorperation, where infrastructure-side sensors supply complementary perception data, effectively mitigating blind spots and occlusions in the vehicle’s field of view. Through the VIC Cross-Attention module, LET-VIC seamlessly integrates multi-view BEV features from infrastructure sensors, bridging information gaps in the single-vehicle perspective. This integration significantly enhances detection accuracy and tracking performance while maintaining robustness, even under delayed conditions. Specifically, in detection performance, LET-VIC achieves substantial improvements over LET-V, with +15.0% mAP at 0ms, +12.3% mAP at 100ms, +9.8% mAP at 200ms and +7.0% mAP at 300ms. For tracking performance, LET-VIC outperforms LET-V with +17.3% AMOTA at 0 ms, +16.4% AMOTA at 100 ms, +14.6% AMOTA at 200 ms and +10.3% AMOTA at 300 ms. Moreover, LET-VIC consistently achieves lower AMOTP values than LET-V, with -0.413 AMOTP at 0 ms, -0.397 AMOTP at 100 ms, -0.338 AMOTP at 200 ms and -0.302 AMOTP at 300 ms, reflecting superior localization accuracy. Additionally, as shown in Fig. 6, although each metric shows a slight decline within a latency range up to 300 ms, LET-VIC still surpasses LET-V by +7.0% in mAP, +10.3% in AMOTA, and -0.302 in AMOTP at 300 ms delay, demonstrating its ability to effectively fuse infrastructure features under certain delays and enhance perception capabilities, showcasing robustness to latency. Furthermore, our experiments on the V2X-Sim dataset [32] validate the generalizability of our approach under more challenging cooperative conditions. Unlike the V2X-Seq-SPD dataset [9], which offers dense collaboration with a 10Hz sampling frequency and full-frame cooperation, V2X-Sim [32] operates at a lower temporal resolution (5Hz) and exhibits sparser cooperation—only 87.2% of frames contain valid collaborative perception due to limited sensor overlap between the RSU and CAV. Despite these constraints, LET-VIC still achieves notable improvements over LET-V, with an increase of +2.2% in mAP and +6.7% in AMOTA. This performance gain under partial cooperation underscores LET-VIC’s capability to effectively leverage intermittent infrastructure signals and maintain robust perception even when continuous collaboration is not feasible. Overall, the consistent performance across these diverse datasets confirms the adaptability of our approach to different sensor configurations and collaboration frequencies.
The comparative results between LET-VIC and models such as V2VNet [18], FFNet [22], and PointPillars [53] indicate that the end-to-end temporal perception approach of LET-VIC significantly outperforms traditional tracking-by-detection models across various delay settings. The performance improvement of LET-VIC is due to its ability to fuse temporal history information. This allows previously visible objects’ features to be carried forward, compensating for occlusions in the current frame. For example, targets visible in prior frames but obscured in the current frame benefit from temporal feature integration, resulting in enhanced detection and tracking performance. This capability is particularly beneficial in addressing blind spots in the current frame by leveraging features from previous observations. Specifically, LET-VIC achieves substantial improvements in detection performance, with at least +13.7% mAP at 0ms, +12.3% mAP at 100ms, +11.0% mAP at 200ms and +8.5% mAP at 300ms compared to the other models. For tracking performance, LET-VIC outperforms V2VNet [18], FFNet [22], and PointPillars [53] with at least +13.1% AMOTA at 0 ms, +12.1% AMOTA at 100 ms, +12.7% AMOTA at 200 ms and +9.2% AMOTA at 300 ms. Moreover, LET-VIC consistently achieves lower AMOTP values, with at least -0.122 AMOTP at 0 ms, -0.117 AMOTP at 100 ms, -0.100 AMOTP at 200 ms and -0.063 AMOTP at 300 ms, reflecting superior localization accuracy. Additionally, Compared to V2VNet [18], FFNet [22], and PointPillars [53], LET-V demonstrates that even with single-vehicle perception, the end-to-end approach significantly improves detection and tracking. Its performance approaches that of PointPillars [53] and surpasses both FFNet [22] and V2VNet [18]. This highlights the effectiveness of the end-to-end paradigm in enhancing perception performance, even without infrastructure data.
IV-C Ablation Study
TABLE III: Importance of CEC in LET-VIC on the V2X-Seq-SPD-2Hz Dataset at Different Latency Levels.
@@ IV-C2 LET-VIC Implementation Details
Due to limited computational resources, this section of experiments extracts the V2X-Seq-SPD dataset [9] at a reduced frequency of 2Hz to generate the V2X-Seq-SPD-2Hz dataset [9]. Additionally, we reduced the feature dimensionality in the model from 256 dimensions to 64 dimensions to expedite training and evaluation. Other settings remain consistent with those described in Section IV-A2.
@@ IV-C3 Analysis
The results presented in Table III demonstrate that LET-VIC with the CEC module consistently outperforms the standard LET-VIC across all delay settings, highlighting the CEC module’s effectiveness in enhancing perception robustness. By compensating for sensor calibration errors, the CEC module improves the accuracy of multi-view feature fusion. This ensures that the LiDAR feature points from the vehicle and infrastructure sides, corresponding to the same real-world location, are correctly aligned. Without such compensation, misalignments between vehicle-side and infrastructure-side feature points may introduce noise during feature fusion, negatively impacting perception performance. Specifically, LET-VIC with calibration error compensation achieves +6.4% mAP and +4.7% AMOTA improvements, along with a -0.038 AMOTP reduction at 0 ms delay; +7.5% mAP, +5.5% AMOTA, and -0.079 AMOTP at 100 ms delay; +5.0% mAP, +1.8% AMOTA, and -0.044 AMOTP at 200 ms delay; and +5.2% mAP, +4.2% AMOTA, and -0.042 AMOTP at 300 ms delay. By compensating for sensor calibration errors, this module improves multi-view feature fusion, leading to better overall performance and demonstrating LET-VIC’s robustness to calibration inaccuracies.